OpenStudy (unklerhaukus):

\[{∂^2 \over ∂x^2} \ln(x^2 + y^2)\] \[{∂^2 \over ∂x^2} \ln(x^2 + y^2)\] @Mathematics

OpenStudy (anonymous):

fxx

OpenStudy (anonymous):

derivate with respect X, treating y or anything else as a constant

OpenStudy (anonymous):

you start with fx, fy, then you have fxx, fxy,fyx,fyy , it turns out that fxy = fyx which is called fubini's theorem

OpenStudy (unklerhaukus):

\[F_x={∂ \over ∂x} ({2x\over x^2 +y^2 })\]

OpenStudy (unklerhaukus):

right so far?

OpenStudy (jamesj):

yes

OpenStudy (unklerhaukus):

cool

OpenStudy (unklerhaukus):

\[= {2 \over x^2 +y^2}+ {-4x^2 \over (x^2+y^2)^2}\]

OpenStudy (unklerhaukus):

\[= {2 (x^2 + y^2) -4x^2 \over (x^2 +y^2)^2}\]

OpenStudy (unklerhaukus):

is this correct ?

OpenStudy (unklerhaukus):

...

OpenStudy (jamesj):

Yes and simplify.

OpenStudy (unklerhaukus):

\[=-{2(x^2-y^2) \over (x^2 + y^2)^2}\]

OpenStudy (jamesj):

Sure, or bring the - inside the numerator.

OpenStudy (unklerhaukus):

yeah i could do that but i am trying to prove that The Laplacian of (ln(x^2 + y^2) =0

OpenStudy (jamesj):

Which should now be obvious as the above result is anti-symmetric in x and y.

OpenStudy (unklerhaukus):

yeah i got it now, Thanks a million JamesJ

OpenStudy (jamesj):

The other way to do it is to use the Laplacian in polar coordinates, as your function is ln(r^2) = 2 ln(r).

OpenStudy (jamesj):

For the record, in polars, \[ \nabla^2 f(r,\theta) = f_{rr} + \frac{1}{r}f_r + \frac{1}{r^2}f_{\theta \theta} \]

OpenStudy (jamesj):

With you function f = 2 ln r, hence \( f_r = 2/r , f_{rr} = -2/r^2 \) and therefore \[ \nabla^2 f = -2/r^2 + (1/r) . (2/r) + 0 = 0 \]

OpenStudy (jamesj):

However, if you don't already have the Laplacian in polars, the Cartesian calculation you're doing is much easier than proving the polar Laplacian! :-)

OpenStudy (unklerhaukus):

i diden't even think of polar coordinates, but that looks like a much more elegant way so solve the question. thanks